OpenStudy (anonymous):

Find the inverse of f a.) 2x/x+7 b.) 4x/3x-7

OpenStudy (anonymous):

@jim_thompson5910 Could you help me with problem please?

jimthompson5910 (jim_thompson5910):

a) is \[\Large f(x) = \frac{2x}{x+7}\] right?

OpenStudy (anonymous):

@jim_thompson5910 it says that's wrong...

jimthompson5910 (jim_thompson5910):

what did you get?

OpenStudy (anonymous):

i didn't get anything cause i didn't know how to approach the problem?

jimthompson5910 (jim_thompson5910):

but you said "it says that's wrong..." that implies you got an answer of some kind

OpenStudy (anonymous):

oh sorry i'm stupid

OpenStudy (anonymous):

yes you were right with \[\frac{ 2x}{ x+7 }\]

jimthompson5910 (jim_thompson5910):

so what you need to do is solve for y \[\Large x=\frac{2y}{y+7}\]

OpenStudy (anonymous):

ok so what do i do from there?

jimthompson5910 (jim_thompson5910):

did you solve for y?

OpenStudy (anonymous):

i didn't know hwo to sorry

jimthompson5910 (jim_thompson5910):

it might help to multiply both sides by y+7 first

OpenStudy (anonymous):

then you'd get... x(x+7)=2y xy+7x=2y

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

why not get all the y terms to one side

jimthompson5910 (jim_thompson5910):

so move the xy over

jimthompson5910 (jim_thompson5910):

and try to isolate y

OpenStudy (anonymous):

wait what? sorry...

jimthompson5910 (jim_thompson5910):

xy+7x=2y 7x=2y-xy ... Subtract xy from both sides.

jimthompson5910 (jim_thompson5910):

the goal is to isolate y

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

that would make it 7x=y(2-x)

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

you have one more step

OpenStudy (anonymous):

divide both sides by y?

OpenStudy (anonymous):

and subtract 7x from both sides?

jimthompson5910 (jim_thompson5910):

if you divided both sides by y, you'd just get y back onto a side with other terms you want to get y all by itself

OpenStudy (anonymous):

how do you do that?

jimthompson5910 (jim_thompson5910):

give it a shot, try to get y all by itself 7x = y*(2-x)

jimthompson5910 (jim_thompson5910):

try to move that 2-x over

OpenStudy (anonymous):

6x+2=y?

OpenStudy (anonymous):

wait that won't work

OpenStudy (anonymous):

the only wait to move the (2-x) is to divide it to both sides right?

jimthompson5910 (jim_thompson5910):

correct

jimthompson5910 (jim_thompson5910):

divide both sides by 2-x

OpenStudy (anonymous):

so the final answer would be 7x/ (2-x)

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

thankyou!! & I tried to do part b by myself this time and I got ... y=-7x/(4-3x) and it says it's wrong, but i don't know what i did wrong

jimthompson5910 (jim_thompson5910):

x = 4y/(3y-7) x(3y-7) = 4y 3xy - 7x = 4y I'll let you finish

OpenStudy (anonymous):

would 3xy-7x=4y would go to be x(3y-7)=4y

OpenStudy (anonymous):

but wiat that wouldn't work

jimthompson5910 (jim_thompson5910):

move the y term over

jimthompson5910 (jim_thompson5910):

get all the y terms together

OpenStudy (anonymous):

so... -4y+3xy-7x=0?

jimthompson5910 (jim_thompson5910):

move the 7x over so the y terms are on their own side

OpenStudy (anonymous):

ok that makes sense

OpenStudy (anonymous):

wait but where would the 3xy be?

jimthompson5910 (jim_thompson5910):

you would have -4y+3xy = 7x

OpenStudy (anonymous):

alright thanks I got it!!

jimthompson5910 (jim_thompson5910):

np